A sum of money lent out at simple interest amounts to 2200in one year and 2800 in 4 years .find the sum of money and rate of interest

Found 2 solutions by Boreal, TeachMath:
Answer by Boreal(15235)

rate is r
x(1+r)=2200
x=2200/(1+r)
we also know
x(1+r)^4=2800
substitute
2200(1+r)^3=2800
(1+r)^3=2800/2200=1.273
1+r=1.084
r=0.084 or 8.4% interest
sum of money is 2200/(1.084)=$2030.06, rounding to end
check by looking for 4 years at 8.4%=$2803, within rounding error.

Answer by TeachMath(96)

correct answer: $2,000 and 10%